OpenStudy (anonymous):

A square pyramid is 6 feet on each side. The height of the pyramid is 4 feet. What is the total area of the pyramid?

OpenStudy (anonymous):

a. 60 ft2 b. 156 ft2 c. 96 ft2 d. 120 ft2

OpenStudy (goldphenoix):

To find the formula of any pyramid, you do: \[\frac{ 1 }{ 3 } (Base x Height)\] A square pyramid is 6 feet on each side. So to find the area of the base, you do 6x6 = 36. Height is 4. 36x4+3= Your answer.

OpenStudy (goldphenoix):

Hm... I did something wrong.

OpenStudy (goldphenoix):

I believe I did the volume, my bad.

OpenStudy (anonymous):

 so the height of the slanted triangle is 5 ft so its area is 1/2*5*6=15 so for 4 triangles it becomes 60 + the area of the base that is 6*6=36ft so total area =36+60=96ft